OpenStudy (anonymous):

change 2 3/5 to tenths

OpenStudy (anonymous):

first make it into one fraction 2 would be 10/5 because its 5/5+5/5 (adding fractions need same dominator) then 10/5+3/5=13/5 then to make fifths into tenths you times by two both bottom and top so its 13/5 X2 which then makes 26/10 which if you want a whole number at the front you look at how many times 26 can go into 10, which is twice and 2X10=20 26-20=6 so your final answer should be 2 6/10 :)